OpenStudy (anonymous):

Write in standard form for an ellipse. 16x^2 + 9y^2 = 1

OpenStudy (anonymous):

16x^2 + 9y^2 = 1 x^2/(1/16) + y^2/(1/9) = 1

OpenStudy (anonymous):

you divided x2 by 1/16 and y2 by 1/9?

OpenStudy (anonymous):

merely rearranged yup

OpenStudy (anonymous):

what would the result look like?

OpenStudy (anonymous):

x^2+y^2=1?

OpenStudy (anonymous):

nope that s the equation of the circle with origin as centre and raius 1 unit

OpenStudy (anonymous):

x^2/(1/16) + y^2/(1/9) = 1 represents ellipse..

OpenStudy (anonymous):

oh so a and b will be fractions

OpenStudy (anonymous):

a=1/9 and b=1/16

OpenStudy (anonymous):

a^2=1/16 and b^2=1/9

OpenStudy (anonymous):

right ok thanks!!! i get it now

OpenStudy (anonymous):

wait but shouldnt a^2=1/9 since its the larger value?

OpenStudy (anonymous):

its up to you to use a^2 or b^2